# InternVideo2.5: Empowering Video MLLMs with Long and Rich Context Modeling

### v1

@@ 1 Introduction
Multimodal large language models (MLLM) make milestones in reaching artificial general intelligence. They integrate different sensing signals into a unified LLM-based framework, reformulating most perception and cognition problems into the multimodal next token prediction task. This successfully addresses applications ranging from multimodal document analysis OpenAI (2024); Reid et al. (2024); Chen et al. (2024c), video understanding Li et al. (2023); Wang et al. (2024e); Li et al. (2024e), agent interactions to scientific discovery Chen et al. (2023a), world modeling Agarwal et al. (2025), autonomous driving Hu et al. (2023), and live assistance Mu et al. (2024); Driess et al. (2023).
@@ 2 Related Work
Multimodal Large Language Models (MLLMs) usually combine vision encoder, LLM, and their connectors through instruction-tuning, capable of handling sophisticated tasks such as generating accurate picture descriptions and responding to visual queries. Progresses have been made to explore its model architecture Zohar et al. (2024); Liu et al. (2025); Li et al. (2023); Wang et al. (2022), size Chen et al. (2024c); Zhang et al. (2024c) and capability Wang et al. (2024c), training corpus Li et al. (2024d); Wang et al. (2023), preference optimization Yan et al. (2024); Yu et al. (2024), and more. Contemporary video-capable MLLMs have demonstrated the ability to process sequential visuals, comprehending spatiotemporal variations. Works discuss different how different vision encoders Zohar et al. (2024) and connectors Liu et al. (2025) affect MLLMs’ performance, and find video encoders are still irreplaceable with image ones. Also, though MLPs related connectors are proven as effective as Q-Former, pooling, and other compression-preferred ones, the latter ones can deliver more efficient solutions, benefiting potential long visuals processing Li et al. (2024e).
@@
While conventional MLLMs have demonstrated strong performance in tasks like visual captioning and question answering, they are challenged in handling fine-grained visual tasks such as segmentation and temporal grounding that require precise predictions. To address these limitations, researchers have developed two main approaches: 1) The pixel-to-sequence (P2S) methodology Chen et al. (2023b); Ren et al. (2024); Wang et al. (2024a, f); Ye et al. (2023) enables MLLMs to generate textual predictions directly. These systems incorporate specialized components like time-aware encoders and sliding video processors to enhance temporal information understanding. 2) The pixel-to-embedding (P2E) approach Bai et al. (2024); Lai et al. (2024); Wang et al. (2024c); Wu et al. (2024b); Zhang et al. (2023a) focuses on compressing visual information before passing it to specialized downstream decoders for final predictions. Recent implementations leverage advanced segmentation tools through prompt-based mechanisms, using special tokens to bridge MLLMs with segmentation modules. Some systems employ multiple routing tokens and enhanced query mechanisms to facilitate connections between MLLMs and various decoder components. These developments represent significant progress in creating more capable and versatile visual understanding systems.
@@ 3.3 Training Video Corpus for Multimodal Context Modeling
We curate a collection of visual-text data with xxx image-text pairs and xxx video-text pairs, as well as some text data for language capability enhancement. For visual-caption pairs, we convert them into a question-answing (QA) format for training convenience. Specifically,
  • Image-Text. We utilize 3.5 million detailed image descriptions recaptioned with LLava-NeXT-34B Zhang et al. (2024c) from COCO118K, BLIP558K, and CC3M datasets Li et al. (2024a). For instruction data, we use single-image instructions from LLava-NeXT Zhang et al. (2024c), Allava Chen et al. (2024a), and ShareGPT4O Chen et al. (2024c); Wang et al. (2024e), along with multi-image instructions from LLaVA-Interleave Li et al. (2024b). Additionally, we incorporate 558K image-text pairs from LCS-558K Liu et al. (2024).
  • Video-Text. Our video-text data consists of WebVid2M Bain et al. (2021) descriptions recaptioned with VideoChat2 Li et al. (2024c), and 323K detailed descriptions from WebVid Bain et al. (2021) and Kinetics Kay et al. (2017) recaptioned with Gemini Reid et al. (2024)Share (2024). For instruction fine-tuning, we use short video data from VideoChat2Li et al. (2024c) and InternVideo2 Wang et al. (2024e), complemented by GPT4o-annotated data from ShareGPT4o Chen et al. (2024c); Wang et al. (2024e), VideoChatGPT-Plus Maaz et al. (2024), LLaVA-Video-178K Zhang et al. (2024d), and LLava-Hound Zhang et al. (2024b).
@@
  • Segmentation. We use DiDeMo Hendricks et al. (2017) and QuerYD Oncescu et al. (2021) for temporal grounding task, RefCOCO Yu et al. (2016), RefCOCOg Yu et al. (2016) and RefCOCO+ Yu et al. (2016) for spatial grounding task, and SAMv2 Ravi et al. (2024), MeViS Ding et al. (2023) for segmentation task.
  • Spatial Grounding. We use AS-V2 Wang et al. (2024b), Visual Genome Krishna et al. (2017), RefCOCO Yu et al. (2016), RefCOCOg Yu et al. (2016), RefCOCO+ Yu et al. (2016) for one epoch with a total batch size of 128 to train region head and token.
@@ 3.5 Implementation
The framework incorporates multiple specialized components: a Temporal Head based on CG-DETR architecture Moon et al. (2023), and a Mask Head utilizing SAM2’s pre-trained weights. For temporal processing, the system leverages InternVideo2 Wang et al. (2024e) for video feature extraction, while query features are processed through the language model. To enhance spatiotemporal capabilities, we implement two-layer MLPs for both positioning prompts and spatial input encoding into the multimodal language model.
@@ 4 Experiments
Model Size #Tokens MVBench PerceptionTest EgoSchema LongVideoBench MLVU VideoMME Average duration (sec) 16 23 180 473 651 1010 Proprietary Models GPT4-V (OpenAI, 2023) - - 43.7 - - 59.1 49.2 59.9 GPT4-o (OpenAI, 2024) - - 64.6 - 72.2 66.7 64.6 71.9 Gemini-1.5-Pro (Reid et al., 2024) - - 60.5 - 71.2 64.0 - 75.0 Open-Source MLLMs InternVL2 (Chen et al., 2024c) 8B 256 66.4 - - - - 54.0 InternVL2 (Chen et al., 2024c) 76B 256 69.6 - - - - 61.2 LLaVA-NeXT-Video (Zhang et al., 2024c) 7B 144 53.1 48.8 - 49.1 - 46.5 LLaVA-OneVision (Li et al., 2024a) 7B 196 56.7 57.1 60.1 56.3 64.7 58.2 LLaVA-OneVision (Li et al., 2024a) 72B 196 59.4 66.9 - 61.3 68.0 66.2 VideoLLaMA2 (Cheng et al., 2024) 7B 72 54.6 51.4 51.7 - 48.5 47.9 VideoLLaMA2 (Cheng et al., 2024) 72B 72 62.0 57.5 63.9 - - 62.4 QwenVL2 (Bai et al., 2023) 7B - 67.0 62.3 66.7 - - 63.3 QwenVL2 (Bai et al., 2023) 72B - 73.6 68.0 77.9 - - 71.2 VideoChat2-HD (Li et al., 2024c) 7B 72 62.3 - - - 47.9 45.3 InternVideo2-HD (Wang et al., 2024e) 7B 72 67.2 63.4 60.0 - - 49.4 VideoChat-TPO Yan et al. (2024) 7B 64 66.8 - - - 54.7 - InternVL2.5 Chen et al. (2024b) 7B 256 72.0 68.2 51.5 60.0 68.9 64.2 Open-Source Long Video MLLMs LLaMA-VID (Li et al., 2024f) 7B 2 41.9 44.6 - - 33.2 25.9 LongVILA (Xue et al., 2024) 7B 196 - - 67.7 - - 57.5 LongVA (Zhang et al., 2024a) 7B 144 - - - - 56.3 52.6 LongLLaVA (Wang et al., 2024d) 9B 144 49.1 - - - - 43.7 LongVU (Shen et al., 2024) 7B 64 66.9 - 67.6 - 65.4 - VideoChat-Flash Li et al. (2024e) 7B 16 73.2 75.6 - 64.2 74.5 63.0 InternVideo2.5 (InternVL2.5+LRC) 7B 16 75.7(+3.7) 74.9(+6.7) 63.9(+12.4) 60.6(+0.6) 72.8(+3.9) 65.1(+0.9)
@@ 4.1 Video Understanding
As given in Table 2, InternVideo2.5 achieves almost the leading performance in all popular short and long video question-answering benchmarks in around 7B LLM capacity level. Compared with the employed base MLLM InternVL2.5, InternVideo2.5 gives a overall increase no matter on short or long duration predictions. The rises are notable when dealing with short video understanding, InternVideo2.5 increases more than 3 points on MVBench and Perception Test based on InternVL2.5. Concerning long video understanding, the whole trend is still rising but the variations are changing on different benchmarks. The brought increases are clear on EgoSchema (+12.4) and MLVU (+3.9) while they are relatively marginal on LongVideoBench (+0.6) and VideoMME (+0.9). We attribute this to the questions in the latter two benchmarks are more dependent of world knowledge and reasoning rather than perceived evidences, since the performance increases over them are more apparent with the larger language models.

### v2

@@ 1 Introduction
Multimodal large language models (MLLM) make milestones in reaching artificial general intelligence. They integrate different sensing signals into a unified LLM-based framework, reformulating most perception and cognition problems into the multimodal next token prediction task. This successfully addresses applications ranging from multimodal document analysis OpenAI (2024); Reid et al. (2024); Chen et al. (2024c), video understanding Li et al. (2023); Wang et al. (2024f); Li et al. (2024e), agent interactions, to scientific discovery Chen et al. (2023a), world modeling Agarwal et al. (2025), autonomous driving Hu et al. (2023), and live assistance Mu et al. (2024); Driess et al. (2023).
@@ 2 Related Work
Multimodal Large Language Models (MLLMs) usually combine vision encoder, LLM, and their connectors through instruction-tuning, capable of handling sophisticated tasks such as generating accurate picture descriptions and responding to visual queries. Progresses have been made to explore its model architecture Zohar et al. (2024); Liu et al. (2025); Li et al. (2023); Wang et al. (2022); Ye et al. (2023, 2024), size Chen et al. (2024c); Zhang et al. (2024c) and capability Wang et al. (2024d), training corpus Li et al. (2024d); Wang et al. (2023); GLM et al. (2024), preference optimization Yan et al. (2024); Yu et al. (2024), and more. Contemporary video-capable MLLMs have demonstrated the ability to process sequential visuals, comprehending spatiotemporal variations. Works discuss different how different vision encoders Zohar et al. (2024) and connectors Liu et al. (2025) affect MLLMs’ performance, and find video encoders are still irreplaceable with image ones. Also, though MLPs related connectors are proven as effective as Q-Former, pooling, and other compression-preferred ones, the latter ones can deliver more efficient solutions, benefiting potential long visuals processing Li et al. (2024e).
@@
While conventional MLLMs have demonstrated strong performance in tasks like visual captioning and question answering, they are challenged in handling fine-grained visual tasks such as segmentation and temporal grounding that require precise predictions. To address these limitations, researchers have developed two main approaches: 1) The pixel-to-sequence (P2S) methodology Chen et al. (2023b); Ren et al. (2024); Wang et al. (2024a, g); Ye et al. (2023) enables MLLMs to generate textual predictions directly. These systems incorporate specialized components like time-aware encoders and sliding video processors to enhance temporal information understanding. 2) The pixel-to-embedding (P2E) approach Bai et al. (2024); Lai et al. (2024); Wang et al. (2024d); Wu et al. (2024b); Zhang et al. (2023a) focuses on compressing visual information before passing it to specialized downstream decoders for final predictions. Recent implementations leverage advanced segmentation tools through prompt-based mechanisms, using special tokens to bridge MLLMs with segmentation modules. Some systems employ multiple routing tokens and enhanced query mechanisms to facilitate connections between MLLMs and various decoder components. These developments represent significant progress in creating more capable and versatile visual understanding systems.
@@ 3.3 Training Video Corpus for Multimodal Context Modeling
We curate a collection of visual-text data with 7M image-text pairs and 3.7M video-text pairs, as well as 143K text data for language capability enhancement. For visual-caption pairs, we convert them into a question-answering (QA) format for training convenience. Specifically,
  • Image-Text. We utilize 3.5 million detailed image descriptions recaptioned with LLava-NeXT-34B Zhang et al. (2024c) from COCO118K, BLIP558K, and CC3M datasets Li et al. (2024a). For instruction data, we use single-image instructions from LLava-NeXT Zhang et al. (2024c), Allava Chen et al. (2024a), and ShareGPT4O Chen et al. (2024c); Wang et al. (2024f), along with multi-image instructions from LLaVA-Interleave Li et al. (2024b). Additionally, we incorporate 558K image-text pairs from LCS-558K Liu et al. (2024).
  • Video-Text. Our video-text data consists of WebVid2M Bain et al. (2021) descriptions recaptioned with VideoChat2 Li et al. (2024c), and 323K detailed descriptions from WebVid Bain et al. (2021) and Kinetics Kay et al. (2017) recaptioned with Gemini Reid et al. (2024)Share (2024). For instruction fine-tuning, we use short video data from VideoChat2Li et al. (2024c) and InternVideo2 Wang et al. (2024f), complemented by GPT4o-annotated data from ShareGPT4o Chen et al. (2024c); Wang et al. (2024f), VideoChatGPT-Plus Maaz et al. (2024), LLaVA-Video-178K Zhang et al. (2024d), and LLava-Hound Zhang et al. (2024b).
@@
  • Segmentation. We use MeViS Ding et al. (2023) and SAMv2 Ravi et al. (2024) for referring segmentation task.
  • Spatial Grounding. We use AS-V2 Wang et al. (2024c), Visual Genome Krishna et al. (2017), RefCOCO Yu et al. (2016), RefCOCOg Yu et al. (2016), RefCOCO+ Yu et al. (2016) for one epoch with a total batch size of 128 to train region head and token.
@@ 3.5 Implementation
The framework incorporates multiple specialized components: a Temporal Head based on CG-DETR architecture Moon et al. (2023), and a Mask Head utilizing SAM2’s pre-trained weights. For temporal processing, the system leverages InternVideo2 Wang et al. (2024f) for video feature extraction, while query features are processed through the language model. To enhance spatiotemporal capabilities, we implement two-layer MLPs for both positioning prompts and spatial input encoding into the multimodal language model.
@@ 4 Experiments
Model Size #Tokens MVBench PerceptionTest EgoSchema LongVideoBench MLVU VideoMME LVBench Average duration (sec) 16 23 180 473 651 1010 4101 Proprietary Models GPT4-V (OpenAI, 2023) - - 43.7 - - 59.1 49.2 59.9 - GPT4-o (OpenAI, 2024) - - 64.6 - 72.2 66.7 64.6 71.9 30.8 Gemini-1.5-Pro (Reid et al., 2024) - - 60.5 - 71.2 64.0 - 75.0 33.1 Open-Source MLLMs InternVL2 (Chen et al., 2024c) 8B 256 66.4 - - - - 54.0 - InternVL2 (Chen et al., 2024c) 76B 256 69.6 - - - - 61.2 - LLaVA-NeXT-Video (Zhang et al., 2024c) 7B 144 53.1 48.8 - 49.1 - 46.5 - LLaVA-OneVision (Li et al., 2024a) 7B 196 56.7 57.1 60.1 56.3 64.7 58.2 - LLaVA-OneVision (Li et al., 2024a) 72B 196 59.4 66.9 - 61.3 68.0 66.2 26.9 VideoLLaMA2 (Cheng et al., 2024) 7B 72 54.6 51.4 51.7 - 48.5 47.9 - VideoLLaMA2 (Cheng et al., 2024) 72B 72 62.0 57.5 63.9 - - 62.4 - mPLUG-Owl3 (Ye et al., 2024) 7B - 54.5 - - 52.1 - 53.5 43.5 QwenVL2 (Bai et al., 2023) 7B - 67.0 62.3 66.7 - - 63.3 - QwenVL2 (Bai et al., 2023) 72B - 73.6 68.0 77.9 - - 71.2 41.3 VideoChat2-HD (Li et al., 2024c) 7B 72 62.3 - - - 47.9 45.3 - InternVideo2-HD (Wang et al., 2024f) 7B 72 67.2 63.4 60.0 - - 49.4 - VideoChat-TPO Yan et al. (2024) 7B 64 66.8 - - - 54.7 - - InternVL2.5 Chen et al. (2024b) 7B 256 72.0 68.2 51.5 60.0 68.9 64.2 38.4 Open-Source Long Video MLLMs LLaMA-VID (Li et al., 2024f) 7B 2 41.9 44.6 - - 33.2 25.9 23.9 LongVILA (Xue et al., 2024) 7B 196 - - 67.7 - - 57.5 - LongVA (Zhang et al., 2024a) 7B 144 - - - - 56.3 52.6 - LongLLaVA (Wang et al., 2024e) 9B 144 49.1 - - - - 43.7 - LongVU (Shen et al., 2024) 7B 64 66.9 - 67.6 - 65.4 - - VideoChat-Flash Li et al. (2024e) 7B 16 73.2 75.6 - 64.2 74.5 64.0 47.2 InternVideo2.5 (InternVL2.5+LRC) 7B 16 75.7(+3.7) 74.9(+6.7) 63.9(+12.4) 60.6(+0.6) 72.8(+3.9) 65.1(+0.9) 46.4(+8.0)
@@ 4.1 Video Understanding
As given in Table 2, InternVideo2.5 achieves almost the leading performance in all popular short and long video question-answering benchmarks in around 7B LLM capacity level. Compared with the employed base MLLM InternVL2.5, InternVideo2.5 gives a overall increase no matter on short or long duration predictions. The rises are notable when dealing with short videos, InternVideo2.5 increases more than 3 points on MVBench and Perception Test based on InternVL2.5. Concerning long video understanding, the whole trend is still rising but the variations are changing on different benchmarks. The brought increases are clear on EgoSchema (test) (+12.4) Mangalam et al. (2023), MLVU (+3.9) Zhou et al. (2024), and LVBench (+8.0) Wang et al. (2024b) while they are relatively marginal on LongVideoBench (+0.6) Wu et al. (2024a) and VideoMME (wo sub) (+0.9) Fu et al. (2024). We attribute this to the questions in the latter two benchmarks are more dependent of world knowledge and reasoning rather than perceived evidences, since the performance increases over them are more apparent with the larger language models.